OpenStudy (anonymous):

Differential Equation SOS !y'X^2=y(x=y)

OpenStudy (anonymous):

I've divided both sides by x

OpenStudy (anonymous):

\[x^2*y'=y(x+y)\]

OpenStudy (anonymous):

and now it's \[x*y'=(\frac{ y }{ x })(x+y)\]

OpenStudy (anonymous):

\[\frac{ y }{ x }=u\]

OpenStudy (anonymous):

and so \[y=u*x\]

OpenStudy (anonymous):

\[y'=U'X+U\]

OpenStudy (anonymous):

@rvc ?

OpenStudy (anonymous):

@somy?

OpenStudy (anonymous):

diff equation?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

@fbi2015 will you be my savior?

OpenStudy (anonymous):

nah?

OpenStudy (anonymous):

@rvc can you be my rainbow?

OpenStudy (anonymous):

lol I promised myself not to sleep before I solve this

rvc (rvc):

oh well i did not get ur question:(

OpenStudy (anonymous):

I need to solve this DE problem \[X^2*y'=y(x+y)\]

OpenStudy (anonymous):

and I tried to solve it (aforementioned)

OpenStudy (anonymous):

any thoughts about it? Am I going the right way?

rvc (rvc):

well just give me a sec let me try

OpenStudy (anonymous):

I'll keep on trying as well

rvc (rvc):

i think this is homogenuos differential equation :)

rvc (rvc):

try \[y=vx\]

rvc (rvc):

@Hipocampus u there?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

I tried that, and it didn't get me any further...

rvc (rvc):

well i got something take y common from the RHS @Hipocampus

OpenStudy (anonymous):

\[\frac{ du }{ u^2 }=\frac{ dx }{ x}\]

OpenStudy (anonymous):

is that what you got?

rvc (rvc):